what is the answer to the permutation 12p10

Found 2 solutions by jim_thompson5910, Edwin McCravy:
Answer by jim_thompson5910(35256)

Use the formula n P r = (n!)/(n-r)!


So 12 P 10 = (12!)/(12-10)! = (12!)/(2!) = 479001600/2 = 239500800


So 12 P 10 = 239500800


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

Answer by Edwin McCravy(20054)



12P10 means:

12 times 11 times 10 times 9 times 8 times 7 times 6 times 5 times 4 times 3.

12×11×10×9×8×7×6×5×4×3

Count them. there are 10 numbers written down starting with 12, then
1 less than 12, then 1 less than that, etc., and ending up with 3.

Now multiply those 10 numbers together with a calculator.

You'll get 239500800

That's the answer.

Edwin